Let X have a uniform distribution on the interval (8, 13). Find the probability that the sum of 2 independent observations of X is greater than 24.
Answer #1

The sum of 2 independent obs of X > 24 happens with what probability?

To get a sum of 24 , we basically have to have the 2 numbers whose sum will be greater than 24

Now, lets see is the chance for that to happen

The range of sum of X1 and X2 ( 2 obs of X) will be from 16 to 26.

P(X1 + X2 > 24)

= (max - (X1+X2)) / (max-min)

= (26-24)/(26-16)

= 2/10

= 0.2